# On automorphism groups of smooth hypersurfaces

### v1

@@ 3. Abelian subgroups and canonical bounds
In this section, we introduce two associated exact sequences (3.1)-(3.2) for any finite linear group using primitive groups and we show an upper bound for abelian subgroups of automorphism groups ${\rm Aut}(F)$ of smooth forms $F$ (Theorem 3.4). Then we introduce the notion of canonical bound for $|{\rm Aut}(F)|$ (Definition 3.5 and Lemma 3.6), which enables us to prove that Fermat forms have the largest automorphism groups among smooth forms with automorphism groups consisting of semi-permutations (Theorem 3.9).
@@ 3.1. Primitive constituents and associated exact sequences
Let $F\in{\mathbb{C}}[x_{1},\ldots,x_{r}]$ be a non-zero homogeneous polynomial of degree $d\geq 3$, where $r\geq 1$. The automorphism group ${\rm Aut}(F)$ is finite but its group structure is in general complicated. In order to study ${\rm Aut}(F)$, we introduce some notions for finite linear groups. First, we recall the notion of primitive linear groups (for more details, see e.g., [Bli17, Col07, Col08a]). Let $\rho:G\rightarrow{\rm GL}(r,{\mathbb{C}})$ be an irreducible linear representation of a finite group $G$. We say $\rho$ is primitive (resp. imprimitive) if ${\mathbb{C}}^{r}$ cannot (resp. can) be decomposed as a direct sum of proper subspaces permuted under the action of $G$. A finite subgroup in ${\rm GL}(r,{\mathbb{C}})$ is called primitive (resp. irreducible) if the underlying representation is primitive (resp. irreducible). A normal abelian subgroup of a finite primitive linear group $G$ is contained in the center $Z(G)$ of $G$ (cf. [Col08a, Lemma 1]). We call a finite subgroup $H\subset{\rm PGL}(r,{\mathbb{C}})$ primitive if $H=\pi(\widetilde{H})$ for some finite primitive subgroup $\widetilde{H}\subset{\rm GL}(r,{\mathbb{C}})$. Since the center $Z(\widetilde{H})\subset\widetilde{H}$ is a subgroup of $Z({\rm GL}(r,{\mathbb{C}}))=\{\lambda I_{r}\mid\lambda\in{\mathbb{C}}^{\times}\}$, we have $H\cong\widetilde{H}/Z(\widetilde{H})$.
Using primitive groups, we introduce two exact sequences associated to finite linear groups. For this purpose and later use, we fix some notations.
Set-up 3.1.
@@
acts primitively (but not necessarily faithfully) on $W_{ij}$. Such permutations form transitive subgroups $K_{i}$ ($i=1,\ldots,m$) of $S_{k_{i}}$, and there is a natural group homomorphism $G\rightarrow K_{i}$. Let $s:=k_{1}+\cdots+k_{m}$. The dimensions $r_{ij}$ of the subspaces $W_{ij}$ are called subdegrees of $G$. Let
@@
We call $H_{ij}$ the primitive constituents of $G$ (of degree $r_{ij}$) belonging to $V_{i}$.
@@
Under $\bf{Set\textrm{-}up}$ 3.1, we call the intersection
@@
the principal subgroup of $G$. Then we call the following exact sequence the first associated exact sequence of $G$
$1\longrightarrow P\longrightarrow G\longrightarrow K_{1}\times\cdots\times K_{m}.$ (3.1)
Let $\widetilde{H}_{ij}:=\{h\mid\,h=g|_{W_{ij}},g\in{\rm Stab}_{G}(W_{ij})\}$. Then we have surjective group homomorphisms $\phi_{ij}:{\rm Stab}_{G}(W_{ij})\longrightarrow\widetilde{H}_{ij}$, $g\mapsto g|W_{ij}$ and $\pi_{ij}:\widetilde{H}_{ij}\longrightarrow H_{ij}$, $h\mapsto[h].$ Thus, $\widetilde{H}_{ij}$ and $H_{ij}$ are primitive subgroups of ${\rm GL}(W_{ij})$ and ${\rm PGL}(W_{ij})$ respectively. Consider the group homomorphism
@@
Let $N:=\ker(\phi)$. Then we call the following exact sequence the second associated exact sequence of $G$
$1\longrightarrow N\longrightarrow P\longrightarrow H_{11}\times\cdots\times H_{1k_{1}}\times\cdots\times H_{mk_{m}}.$ (3.2)
Clearly, a finite linear group determines uniquely its first and second associated exact sequences up to isomorphisms of exact sequences. After choosing basis for all $W_{ij}$, we obtain a basis of ${\mathbb{C}}^{r}$ and we may identity ${\rm GL}(W_{ij})$ (resp. ${\rm PGL}(W_{ij})$) with ${\rm GL}(r_{ij},{\mathbb{C}})$ (resp. ${\rm PGL}(r_{ij},{\mathbb{C}})$). Then each element $A$ in $P$ can be viewed as a block diagonal matrix $A={\rm diag}(A_{11},\ldots,A_{mk_{m}})$ with blocks $A_{ij}\in{\rm GL}(r_{ij},{\mathbb{C}})$. Note that $N$ is abelian and its elements are of the form ${\rm diag}(\lambda_{11}I_{r_{11}},\lambda_{12}I_{r_{12}},\ldots,\lambda_{mk_{m}}I_{r_{mk_{m}}})$, where $\lambda_{ij}$ are non-zero complex numbers.
@@
Under $\bf{Set\textrm{-}up}$ 3.1, let $l(G):=(r_{1}^{\prime},\ldots,r_{s}^{\prime})$ be the $s$-tuple of positive integers such that $r_{1}^{\prime}\geq r_{2}^{\prime}\geq\cdots\geq r_{s}^{\prime}$ and there exists a permutation $\sigma\in S_{s}$ satisfying $(r_{\sigma(1)}^{\prime},\ldots,r_{\sigma(s)}^{\prime})=(r_{11},r_{12},\ldots,r_{mk_{m}})$. We call $l(G)$ the subdegree sequence of $G$ (of length $s$) and the $s$-tuple $(H_{11},H_{12},\ldots,H_{m(k_{m}-1)},H_{mk_{m}})$ a primitive-constituent sequence of $G$. The $m$-tuple of integers $(k_{1},\dots,k_{m})$ is called an intrinsic multiplicity sequence of $G$.
@@ 3.2. Abelian subgroups and canonical bounds
The following bound for abelian subgroups of the automorphism groups of smooth forms plays a key role in our study.
@@
Let $F=F(x_{1},\dots,x_{r})$ be a smooth homogeneous polynomial of degree $d\geq 3$, where $r\geq 2$. Fix positive integers $r_{1},\dots,r_{m}$, where $1\leq m\leq r$. Let $N$ be an abelian subgroup of ${\rm Aut}(F)$. Suppose $N$ is generated by matrices of the form ${\rm diag}(\lambda_{1}I_{r_{1}},\lambda_{2}I_{r_{2}},\dots,\lambda_{m}I_{r_{m}})$. Then there exists an integer $k\geq 1$ and a partition $m=b+\sum_{i=1}^{k}a_{i}$ for some integers $b\geq 0$, $a_{i}>0$ ($i=1,\ldots,k$) such that $|N|$ divides $(d-1)^{b}\prod_{i=1}^{k}((d-1)^{a_{i}}-(-1)^{a_{i}})$. In particular, $|N|\leq d^{m}$.
@@
We set $s_{0}:=0$ and $s_{i}:=\sum_{j=1}^{i}r_{j}$ for $i=1,\dots,m$. Let $M=\{1,2,\dots,m\}$ and
$\mathcal{P}:=\{(i,j)\in M\times M\mid\;x_{s_{i}}^{d-1}x_{k_{j}}\in F\text{ forsome }k_{j}\;\text{ satisfying }s_{j-1}<k_{j}\leq s_{j}\}.$
By smoothness of $F$ and [OY19, Proposition 3.3], the projection map $p_{1}:\mathcal{P}\rightarrow M$ given by $(i,j)\mapsto i$ must be surjective. Consider the subgroup
$G_{\mathcal{P}}^{(m)}:=\{(\lambda_{1},\lambda_{2},\dots,\lambda_{m})|\;\lambda_{i}^{d-1}\lambda_{j}=1\text{ if }(i,j)\in\mathcal{P}\}\subset{{\mathbb{C}}^{\times}}\times\cdots\times{\mathbb{C}}^{\times}.$
Let $A={\rm diag}(\lambda_{1}I_{r_{1}},\lambda_{2}I_{r_{2}},\dots,\lambda_{m}I_{r_{m}})\in N$. Since $N$ is a subgroup of ${\rm Aut}(F)$, we have $A(F)=F$, which implies that $\lambda_{i}^{d-1}\lambda_{j}=1$ if $x_{s_{i}}^{d-1}x_{k_{j}}\in F$ and $s_{j-1}<k_{j}\leq s_{j}$. Then $N$ is isomorphic to a subgroup of $G_{\mathcal{P}}^{(m)}$ by the well-defined injective morphism $N\rightarrow G_{\mathcal{P}}^{(m)}$, $A\mapsto(\lambda_{1},\dots,\lambda_{m})$.
We may choose a subset $\mathcal{P}^{\prime}$ of $\mathcal{P}$ such that the cardinality of $\mathcal{P}^{\prime}$ is $m$ and $p_{1}(\mathcal{P}^{\prime})=M$. Note that $G_{\mathcal{P}}^{(m)}$ is a subgroup of $G_{\mathcal{P}^{\prime}}^{(m)}$. We claim that there exists an integer $k\geq 1$ and a partition $m=b+\sum_{i=1}^{k}a_{i}$ for some integers $b\geq 0$, $a_{i}>0$ ($i=1,\ldots,k$) such that
$|G_{\mathcal{P}^{\prime}}^{(m)}|=(d-1)^{b}\prod_{i=1}^{k}((d-1)^{a_{i}}-(-1)^{a_{i}}).$
Since $|N|$ divides $|G_{\mathcal{P}^{\prime}}^{(m)}|$ and the value on the right hand side is clearly at most $d^{m}$, it suffices to prove the claim. Let $p_{2}:\mathcal{P}^{\prime}\rightarrow M$ be the projection map given by $(i,j)\mapsto j$. There are two cases: (1) $p_{2}$ is surjective; (2) $p_{2}$ is not surjective.
For case (1), let $k$ be the largest integer such that $M$ can be expressed as a disjoint union of $k$ subsets $M_{i}$ ($i=1\dots,k$) with
$\mathcal{P}^{\prime}=\bigcup_{i=1}^{k}\mathcal{P}^{\prime}\cap(M_{i}\times M_{i}).$
We set $\mathcal{P}^{\prime}_{i}:=\mathcal{P}^{\prime}\cap(M_{i}\times M_{i})$. Then the cardinalities of $\mathcal{P}^{\prime}_{i}$ and $M_{i}$ are equal and we have $G_{\mathcal{P}^{\prime}}^{(m)}\cong G_{\mathcal{P}^{\prime}_{1}}^{(a_{1})}\times\cdots\times G_{\mathcal{P}^{\prime}_{k}}^{(a_{k})}$, where $a_{i}$ is the cardinality of the set $M_{i}$. Note that $m=\sum_{i=1}^{k}a_{i}$. By direct calculation, $|G_{\mathcal{P}^{\prime}_{i}}^{(a_{i})}|=(d-1)^{a_{i}}-(-1)^{a_{i}}$ and $|G_{\mathcal{P}^{\prime}}^{(m)}|=\prod_{i=1}^{k}((d-1)^{a_{i}}-(-1)^{a_{i}})$.
For case (2), we may assume $m\notin p_{2}(\mathcal{P}^{\prime})$. If $m=2$, then $G_{\mathcal{P}^{\prime}}^{(2)}=\{(\lambda_{1},\lambda_{2})|\;\lambda_{1}^{d}=\lambda_{2}^{d-1}\lambda_{1}=1\}$ of order $d(d-1)$. To prove the case $m\geq 3$ inductively, we set $\mathcal{P}^{\prime\prime}:=\mathcal{P}^{\prime}\cap(M^{\prime}\times M^{\prime})$, where $M^{\prime}=M\setminus\{m\}$. Then we have
$\mathcal{P}^{\prime}=\{(m,j_{m})\}\cup\mathcal{P}^{\prime\prime},$
where $j_{m}\in M^{\prime}$. The claim holds for $G_{\mathcal{P}^{\prime\prime}}^{(m-1)}$ by either case (1) or the induction hypothesis. Clearly, there is a short exact sequence
$1\rightarrow C_{d-1}\rightarrow G_{\mathcal{P}^{\prime}}^{(m)}\rightarrow G_{\mathcal{P}^{\prime\prime}}^{(m-1)}\rightarrow 1,$
where the third morphism is given by $(\lambda_{1},\dots,\lambda_{m})\mapsto(\lambda_{1},\dots,\lambda_{m-1})$. Thus $|G_{\mathcal{P}^{\prime}}^{(m)}|=(d-1)\cdot|G_{\mathcal{P}^{\prime\prime}}^{(m-1)}|$, which implies the claim for $G_{\mathcal{P}^{\prime}}^{(m)}$ since the claim holds for $G_{\mathcal{P}^{\prime\prime}}^{(m-1)}$. This completes the proof of the claim and hence the theorem. ∎
Now we are ready to introduce a bound for groups of automorphisms of smooth forms.
@@
$B(G,F)=d^{s}\prod_{i=1}^{s}|H_{i}|\prod_{j=1}^{m}{k_{j}!}.$
@@
Following [OY19, Example 2.1 (17)], let $F=((x_{1}^{4}+x_{2}^{4})+(2+4\xi_{3}^{2})x_{1}^{2}x_{2}^{2})x_{3}+(-(x_{1}^{4}+x_{2}^{4})+(2+4\xi_{3}^{2})x_{1}^{2}x_{2}^{2})x_{4}+x_{3}^{4}x_{4}+x_{4}^{4}x_{3}+x_{5}^{5}$ and the subgroup $\widetilde{G}$ of ${\rm Aut}(F)$
@@
We denote by $(e_{1},e_{2},e_{3},e_{4},e_{5})$ the underlying basis of ${\mathbb{C}}^{5}$. Then using notations in $\bf{Set\textrm{-}up}$ 3.1, we have ${\mathbb{C}}^{5}=V_{1}\oplus V_{2}\oplus V_{3}$, $V_{1}=W_{11}$, $V_{2}=W_{21}\oplus W_{22}$, $V_{3}=W_{31}$, where $V_{1}=\langle e_{1},e_{2}\rangle$, $V_{2}=\langle e_{3},e_{4}\rangle$, $V_{3}=\langle e_{5}\rangle$, $W_{21}=\langle e_{3}\rangle$, $W_{22}=\langle e_{4}\rangle$. The linear group $\widetilde{G}$ has $4$ primitive constituents $H_{ij}$. Moreover, $S_{4}\cong H_{11}\subset{\rm PGL}(2,{\mathbb{C}})$, $|H_{21}|=|H_{22}|=|H_{31}|=1$, the subdegree sequence $l(\widetilde{G})=(2,1,1,1)$, and $\widetilde{G}$ has an intrinsic multiplicity sequence $(1,2,1)$. Then we have $B(\widetilde{G},F)=5^{4}\cdot 24\cdot 2=30000>480=|\widetilde{G}|$, as predicted by Lemma 3.6.
The following simple fact is useful for our study.
Lemma 3.8.
We say an invertible matrix $A$ is semi-permutation if $A$ is a diagonal matrix up to permutation of columns. Note that up to linear change of coordinates, a finite linear group consists of semi-permutation matrices if and only if its subdegree sequence is $(1,\ldots,1)$. As an immediate consequence of Lemma 3.6, we obtain an optimal bound for semi-permutation automorphism groups of smooth hypersurfaces.
@@
Since ${\rm Aut}(F)$ has the subdegree sequence $(1,\ldots,1)$ of length $s=r$ and the intrinsic multiplicity sequence $(k_{1},\ldots,k_{m})$ with $k_{1}+\cdots+k_{m}=r$, by Lemmas 3.6 and 3.8, we have
$|{\rm Aut}(F)|\leq B({\rm Aut}(F),F)=d^{r}\prod_{j=1}^{m}{k_{j}!}\leq d^{r}\cdot r!.$
If $|{\rm Aut}(F)|=d^{r}\cdot r!$, then $N\cong C_{d}^{r}$, $m=1$, and $K_{1}\cong S_{r}$ (here $N$ and $K_{1}$ are as in $\mathbf{Set{\textrm{-}}up}$ 3.1 and Definition 3.2). In this case, $\{x_{i}^{d}\}_{1\leq i\leq r}$ are the only monomials preserved by $N$, thus $F$ is a Fermat form, up to replacing $x_{i}$ by their non-zero multiples. ∎

### v2

@@ 3. Abelian subgroups and canonical bounds
In this section, we introduce two associated exact sequences (3.1)-(3.2) for any finite linear group using primitive groups in §3.1. In §3.2, we show an upper bound for the size of abelian subgroups of automorphism groups ${\rm Aut}(F)$ of smooth forms $F$ (Theorem 3.5), introduce the notion of canonical bound for $|{\rm Aut}(F)|$ (Definition 3.6 and Lemma 3.7), and prove that Fermat forms have the largest automorphism groups among smooth forms with automorphism groups consisting of semi-permutations (Theorem 3.9).
@@ 3.1. Primitive constituents and associated exact sequences
First, we recall the notion of primitive linear groups (for more details, see e.g., [Bli17, Col07, Col08a]).
Definition 3.1.
Let $\rho:G\rightarrow{\rm GL}(r,{\mathbb{C}})$ be an irreducible linear representation of a finite group $G$. We say $\rho$ is primitive (resp. imprimitive) if ${\mathbb{C}}^{r}$ cannot (resp. can) be decomposed as a direct sum of proper subspaces permuted under the action of $G$. A finite subgroup in ${\rm GL}(r,{\mathbb{C}})$ is called primitive (resp. irreducible) if the underlying representation is primitive (resp. irreducible). We call a finite subgroup $H\subset{\rm PGL}(r,{\mathbb{C}})$ primitive if $H=\pi(\widetilde{H})$ for some finite primitive subgroup $\widetilde{H}\subset{\rm GL}(r,{\mathbb{C}})$.
A normal abelian subgroup of a finite primitive linear group $G$ is contained in the center $Z(G)$ of $G$ (see [Col08a, Lemma 1]). Let $\widetilde{H}$ be a finite primitive subgroup in ${\rm GL}(r,{\mathbb{C}})$ and $H:=\pi(\widetilde{H})\subset{\rm PGL}(r,{\mathbb{C}})$. Since the center $Z(\widetilde{H})\subset\widetilde{H}$ is a subgroup of $Z({\rm GL}(r,{\mathbb{C}}))=\{\lambda I_{r}\mid\lambda\in{\mathbb{C}}^{\times}\}$, we have $H\cong\widetilde{H}/Z(\widetilde{H})$. Using primitive groups, we introduce two exact sequences associated to finite linear groups. For this purpose and later use, we fix some notations.
Set-up 3.2.
@@
acts primitively (but not necessarily faithfully) on $W_{ij}$ (see [Col07, Lemma 1]). Such permutations form transitive subgroups $K_{i}$ ($i=1,\ldots,m$) of $S_{k_{i}}$, and there is a natural group homomorphism $\psi_{i}:G\rightarrow K_{i}$. Let $s:=k_{1}+\cdots+k_{m}$. The dimensions $r_{ij}$ of the subspaces $W_{ij}$ are called subdegrees of $G$. Let
@@
We call the $H_{ij}$’s the primitive constituents of $G$ (of degree $r_{ij}$) belonging to $V_{i}$.
@@
Under $\bf{Set\textrm{-}up}$ 3.2, we call the intersection
@@
the principal subgroup of $G$. Consider the group homomorphism
$\psi:G\longrightarrow K_{1}\times\cdots\times K_{m},\;\;g\mapsto(\psi_{1}(g),\ldots,\psi_{m}(g)).$
Note that ${\rm Ker}(\psi)=P$. Then we call the following exact sequence the first associated exact sequence of $G$
[…]
Let $\widetilde{H}_{ij}:=\{h\mid\,h=g|_{W_{ij}},g\in{\rm Stab}_{G}(W_{ij})\}$. Then we have surjective group homomorphisms $\phi_{ij}:{\rm Stab}_{G}(W_{ij})\rightarrow\widetilde{H}_{ij}$, $g\mapsto g|W_{ij}$ and $\pi_{ij}:\widetilde{H}_{ij}\rightarrow H_{ij}$, $h\mapsto[h].$ Thus, $\widetilde{H}_{ij}$ and $H_{ij}$ are primitive subgroups of ${\rm GL}(W_{ij})$ and ${\rm PGL}(W_{ij})$ respectively. Consider the group homomorphism
@@
Let $N:={\rm Ker}(\phi)$. Then we call the following exact sequence the second associated exact sequence of $G$
$1\longrightarrow N\longrightarrow P\xrightarrow{\,\;\phi\;\,}H_{11}\times\cdots\times H_{1k_{1}}\times\cdots\times H_{mk_{m}}.$ (3.2)
Clearly, the decomposition of the $V_{i}$’s as in $\bf{Set\textrm{-}up}$ 3.2 guarantees the existence of the first and second associated exact sequences, which is a key point of our approach. After choosing a basis for each $W_{ij}$, we obtain a basis of ${\mathbb{C}}^{r}$ and we may identify ${\rm GL}(W_{ij})$ (resp. ${\rm PGL}(W_{ij})$) with ${\rm GL}(r_{ij},{\mathbb{C}})$ (resp. ${\rm PGL}(r_{ij},{\mathbb{C}})$). Then each element $A$ in $P$ can be viewed as a block diagonal matrix $A={\rm diag}(A_{11},\ldots,A_{mk_{m}})$ with blocks $A_{ij}\in{\rm GL}(r_{ij},{\mathbb{C}})$. Note that $N$ is abelian and its elements are of the form ${\rm diag}(\lambda_{11}I_{r_{11}},\lambda_{12}I_{r_{12}},\ldots,$ $\lambda_{mk_{m}}I_{r_{mk_{m}}})$, where $\lambda_{ij}$ are non-zero complex numbers.
@@
Under $\bf{Set\textrm{-}up}$ 3.2, let $l(G):=(r_{1}^{\prime},\ldots,r_{s}^{\prime})$ be the $s$-tuple of positive integers such that $r_{1}^{\prime}\geq r_{2}^{\prime}\geq\cdots\geq r_{s}^{\prime}$ and there exists a permutation $\sigma\in S_{s}$ satisfying $(r_{\sigma(1)}^{\prime},\ldots,r_{\sigma(s)}^{\prime})=(r_{11},r_{12},\ldots,r_{mk_{m}})$. We call $l(G)$ the subdegree sequence of $G$ (of length $s$) and the $s$-tuple $(H_{11},H_{12},\ldots,H_{m(k_{m}-1)},H_{mk_{m}})$ a primitive-constituent sequence of $G$. The $m$-tuple of integers $(k_{1},\dots,k_{m})$ is called an intrinsic multiplicity sequence of $G$.
Throughout the paper, for any finite subgroup of ${\rm GL}(r,{\mathbb{C}})$, we fix a primitive-constituent sequence once and for all.
@@ 3.2. Abelian subgroups and canonical bounds
It is known that for a smooth homogeneous polynomial $F(x_{1},\dots,x_{r})$ of degree $d\geq 3$, every abelian subgroup of ${\rm Aut}(F)$ has order at most $d^{r}$ (see [HS80, Lemma 3.1], [Sza96, Abelian Lemma]). The following more refined bound for abelian subgroups of the automorphism groups of smooth forms plays a key role in our study.
@@
Let $F=F(x_{1},\dots,x_{r})$ be a smooth homogeneous polynomial of degree $d\geq 3$, where $r\geq 2$. Fix positive integers $r_{1},\dots,r_{m}$ with $\sum_{1\leq i\leq m}{r_{i}}=r$, where $1\leq m\leq r$. Let $N\subset{\rm Aut}(F)$ be an abelian subgroup generated by matrices of the form ${\rm diag}(\lambda_{1}I_{r_{1}},\lambda_{2}I_{r_{2}},\dots,\lambda_{m}I_{r_{m}})$. Then $|N|\leq d^{m}$.
@@
We prove the theorem by adapting the proof of [Sza96, Bézout Lemma]. If $r=2$, then $|N|\leq d^{m}$ by direct calculation. If $r\geq 3$, then $X:=X_{F}\subset{\mathbb{P}}^{r-1}$ is a smooth hypersurface of degree $d$. Fix $x=(a_{1}:\cdots:a_{r})\in X$ with $a_{i}\neq 0$ for all $i$. Let $T\subset{\rm PGL}(r,{\mathbb{C}})$ be the (abelian) subgroup generated by all invertible diagonal matrices. The stabilizer $T_{x}$ of $x$ is trivial and we identify $T$ with the orbit $Tx$ of $x$ via the $T$ action. We set $U=X\cap T$ and we have ${\rm Lin}(X)\cap T=\bigcap_{s\in U}\,Us^{-1}$. Let $H\subset T$ be the subgroup generated by all matrices of the form ${\rm diag}(\lambda^{\prime}_{1}I_{r_{1}},\lambda^{\prime}_{2}I_{r_{2}},\dots,\lambda^{\prime}_{m}I_{r_{m}})$, where $\lambda^{\prime}_{i}\in{\mathbb{C}}^{\times}$ ($1\leq i\leq m$). Let $Y\subset{\mathbb{P}}^{r-2}$ denote the closure of the orbit $Hx$ of $x$. Note that $Y\cong{\mathbb{P}}^{m-1}$. Then ${\rm Lin}(X)\cap H$=${\rm Lin}(X)\cap T\cap H$=$\bigcap_{s\in U}\,(Us^{-1}\cap H)$. The cardinality of $\bigcap_{s\in U}\,(Us^{-1}\cap H)$ is at most that of $\bigcap_{s\in U}\,(Us^{-1}\cap Y)$. The finiteness of $\bigcap_{s\in U}\,(Us^{-1}\cap Y)$ implies that for $m-1$ general points $s_{0},s_{1},...,s_{m-2}$, the intersection $\bigcap_{i=0}^{m-2}\,(Us_{i}^{-1}\cap Y)$ is already finite. The cardinality of this intersection is at most ${\rm deg}(\bigcap_{i=0}^{m-2}(Xs_{i}^{-1}\cap Y))$. By applying Bézout’s Theorem ([Ful98, Example 8.4.6]), we obtain $|{\rm Lin}(X)\cap H|\leq d^{m-1}$. Then $|N|\leq d\cdot|\pi(N)|\leq d^{m}$ since $\pi(N)\subseteq({\rm Lin}(X)\cap H)$.∎
The proof, using the ideas in [CS95] and [Sza96], is suggested by a referee. A longer proof of Theorem 3.5 without using Bézout’s Theorem can be found in [YYZ24, Theorem 3.4]. Now we are ready to introduce a bound for the size of groups of automorphisms of smooth forms.
@@
$B(G,F)=d^{s}\cdot\prod_{i=1}^{s}|H_{i}|\cdot\prod_{j=1}^{m}{k_{j}!}.$
@@
Following [OY19, Example 2.1 (17)], let $F=\left(\left(x_{1}^{4}+x_{2}^{4}\right)+\left(2+4\xi_{3}^{2}\right)x_{1}^{2}x_{2}^{2}\right)x_{3}+\left(-\left(x_{1}^{4}+x_{2}^{4}\right)+\left(2+4\xi_{3}^{2}\right)x_{1}^{2}x_{2}^{2}\right)x_{4}+x_{3}^{4}x_{4}+x_{4}^{4}x_{3}+x_{5}^{5}$ and the subgroup $\widetilde{G}$ of ${\rm Aut}(F)$
@@
We denote by $(e_{1},e_{2},e_{3},e_{4},e_{5})$ the of ${\mathbb{C}}^{5}$. Then using notations in $\bf{Set\textrm{-}up}$ 3.2, we have ${\mathbb{C}}^{5}=V_{1}\oplus V_{2}\oplus V_{3}$, $V_{1}=W_{11}$, $V_{2}=W_{21}\oplus W_{22}$, $V_{3}=W_{31}$, where $V_{1}=\langle e_{1},e_{2}\rangle$, $V_{2}=\langle e_{3},e_{4}\rangle$, $V_{3}=\langle e_{5}\rangle$, $W_{21}=\langle e_{3}\rangle$, $W_{22}=\langle e_{4}\rangle$. The linear group $\widetilde{G}$ has $4$ primitive constituents $H_{ij}$. Moreover, $S_{4}\cong H_{11}\subset{\rm PGL}(2,{\mathbb{C}})$, $|H_{21}|=|H_{22}|=|H_{31}|=1$, the subdegree sequence $l(\widetilde{G})=(2,1,1,1)$, and $\widetilde{G}$ has an intrinsic multiplicity sequence $(1,2,1)$. Then we have $B(\widetilde{G},F)=5^{4}\cdot 24\cdot 2=30000>480=|\widetilde{G}|$, as predicted by Lemma 3.7.
We say an invertible matrix $A$ is semi-permutation if $A$ is a diagonal matrix up to permutation of columns. Note that up to linear change of coordinates, a finite linear group consists of semi-permutation matrices if its subdegree sequence is $(1,\dots,1)$. Using Theorem 3.5, we obtain an optimal bound for the size of semi-permutation automorphism groups of smooth forms.
@@
We identify $S_{r}$ with the group of all permutations of the coordinates $x_{1},...,x_{r}$. Let $N_{1}\subseteq{\rm Aut}(F)$ be the (abelian) subgroup consisting of all diagonal matrices in ${\rm Aut}(F)$. Since ${\rm Aut}(F)$ only has semi-permutations, $N_{1}$ is normal in ${\rm Aut}(F)$ and the quotient group $Q_{1}:={\rm Aut}(F)/N_{1}$ is isomorphic to a subgroup of $S_{r}$. Then by Theorem 3.5, we have $|{\rm Aut}(F)|\leq d^{r}\cdot r!$ and the equality holds only if $|N_{1}|=d^{r}$ and $Q_{1}\cong S_{r}$. We assume now $|{\rm Aut}(F)|=d^{r}\cdot r!$. If $x_{1}^{d}\in F$, then $x_{i}^{d}\in F$ for all $i\in\{1,2,...,r\}$ by $Q_{1}\cong S_{r}$. Thus, for every $A={\rm diag}(\lambda_{1},...,\lambda_{r})\in N_{1}$, we have $\lambda_{i}^{d}=1$ for all $i$ by $A(F)=F$. Then $N_{1}=\{{\rm diag}(\lambda_{1},...,\lambda_{r})|\,\lambda_{1}^{d}=\cdots=\lambda_{r}^{d}=1\},$ which implies that $\{x_{i}^{d}\}_{1\leq i\leq r}$ are the only monomials preserved by $N_{1}$ and $F$ is a Fermat form, up to replacing $x_{i}$ by their non-zero multiples. If $x_{1}^{d}\notin F$, then by [OY19, Proposition 3.3] and $Q_{1}\cong S_{r}$, we have $x_{i}^{d-1}x_{j}\in F$ for all $i,j$ with $i\neq j$. Then $N_{1}\subseteq N_{2}:=\{{\rm diag}(\lambda_{1},...,\lambda_{r})|\,\lambda_{i}^{d-1}\lambda_{j}=1\text{ for all $i,j$ with $i\neq j$}\},$ which is a contradiction since $d^{r}=|N_{1}|\leq|N_{2}|<d^{r}$ by direct computation. ∎